OpenStudy (anonymous):

What are the real zeroes of x3 + 6 x2 – 9x – 54?

OpenStudy (anonymous):

where the graph is cutting the x axis i.e. y=0

OpenStudy (anonymous):

A. 1, 2, 27B. 3, -3, -6C. -6, 3, -6D. 2, -1, 18E. 3, 3, -6

OpenStudy (anonymous):

Factor by grouping: x^2(x + 6) - 9(x+6) (x^2 - 9)(x+ 6) = 0 so x^2 - 9 = 0 and x + 6 = 0 You can now find the real values of x.

OpenStudy (anonymous):

x^2 - 9 = 0 (x-3)(x+3) = 0 x = 3 and x = -3 x + 6 = 0 x = -6 So the real zeros are 3,-3, and -6

OpenStudy (anonymous):

An entrepreneur estimates that the total profit (profit = total revenue minus total cost) from his proposed company will be given by the function P(x) = x3 – 4x2 + 3x – 12, where P is in hundreds of dollars and x is the number of years elapsed after the company starts operations. In how many years (x) will the company break even (no profit, no loss)? A. 1 year B. 2 years C. 3 years D. 4 years E. 6 years

OpenStudy (anonymous):

How many negative roots can this polynomial function possibly have? f(x) = 5x4 – 7x3 – x2 – 3x – 2 A. 1 B. 5 or 3 or 1 C. 3 or 1 D. 4 or 2 or 0 E. 2 or 0

OpenStudy (anonymous):

What is the remainder of (6x3) (x + 2)?

OpenStudy (anonymous):

Using synthetic division, find (x6 − 250x2 − 96) (x − 4).

OpenStudy (agent0smith):

Maybe one question at a time :P